4. An exam paper consists of ten multiple choice questions, each offering four choices of which only one is correct. If a candidate chooses his answers I completely at random, what is the probability that (i) he gets at least 8 questions right, (ii) the last of the ten questions is the eighth one he gets right, (ii) in six such exams, he gets at least 8 questions right in at most one exam?
Answer #1

The probability of getting 1 question right is:

p = 0.25

This because each question has 4 choices of which only one is correct.

1)  The probability of getting at least 8 questions right of 10 questions is:

P(Xgeq 8)= P(X=8)+P(X=9)+P(X=10)

ー10Cs × 0.258 × 0.752 + 10C, x 0.259 × 0.751 + 10C10 × 0.2510

0.000416

2)  In this part we have to calculate the probability of getting 7 questions right out of 9 questions and 10th questions right. That will make the 10th question the 8th one right.

P = P(X = 7) × 0.25

  C × 0.25 × 0.75° × 0.25 = 0.000309

3)  The probability of getting at least 8 questions right in 1 exam is calculated in part 1. Now the probability of getting at least 8 questions right in atmost 1 exam of 6 exams is:

PlX < 1)一6C1 × 0.000416| × (1-0.000416)5 + 6C0 × (1-0.000416)

-0,999997